Dozier Company produced and sold 1000 units during ts first month of operations, it reported the following costs and expenses for the month: Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead Total manufacturing overhead Variable selling expense Fixed selling expense Total selling expense Variable administrative expense Fixed administrative expense Total administrative expense 71,000 36,000 44,400 - % 31,200 2 29,600 % 15,800 28,600 12,400 18,800 s 4,200 25,400 Required: 1. With respect to cost classifications for preparing financlal statements: a. What is the total product cost? b. What is the total period cost? 3 of 6 Next>
media%2F4fb%2F4fb6035b-7238-4be9-b997-05
Answer #1

Answer(1): a- Product cost = Direct material + Direct labor + Total manufacturing Overheads

= 71000 + 36000 + 44400

Product cost = $151400

b- Period cost = Total selling exps. + Total administrative exps.

= 31200 + 29600

Period cost = $60800

Answer(2): a- Direct manufacturing cost = Direct material + Direct labor

Direct manufacturing cost: 71000 + 36000 = $107000

b- Indirect manufacturing cost = Manufacturing overheads

Indirect manufacturing cost = $44400

Answer(3): Total manufacturing cost = Direct material + Direct labor + Total manufacturing Overheads

= 71000 + 36000 + 44400

= $151400

Total non-manufacturing cost = Total selling exps. + Total administrative exps.

: 31200 + 29600 = $60800

Answer(4): a- Variable manufacturing cost = $15800 (variable manufacturing overheads)

b- Total fixed cost of company as a whole = Fixed manufacturing overhead + Fixed selling exps.+Fixed administrative exps.

Total fixed cost of company as a whole: 28600+18800+25400 = $72800

c- Variable cost per unit produced: 15800 / 1000 = $15.80

Variable cost per unit sold: (12400+4200)/1000 = $16.60